# Machine Learning Framework for Characterizing Processing–Structure Relationship in Block Copolymer Thin Films

### v1

@@ 3.3 Processing-structure relationship prediction
While the SHAP evaluations for the GISAXS-measured FWHM and domain spacing models share some similarities, key differences also emerge. Interestingly, the additive type has a much more significant mean SHAP value for the FWHM predictions, indicating the additive type has a more profound impact on BCP FWHM than domain spacing (Fig. 10 a left). However, additive ratio still maintains the largest mean SHAP value, 0.00029, followed by additive type, 0.00018, and solvent ratio, 0.000054. Similar to the local SHAP analysis for domain spacing, high additive ratios cluster at lower SHAP values, reflecting rapid ordering of the BCP, whereas lower additive ratios exhibit larger and more scattered SHAP values, indicative of weakly-ordered samples with broad FWHM. (Fig. 10 a right). Although the SHAP values for additive type show a similar trend for those related to the domain spacing model, they are far more distinctly separated in the FWHM model, indicating substantial variation in FWHM with changes of additive type. Diving deeper into the SHAP analysis for the FWHM prediction, an interesting relationship can be observed between solvent ratio and additive ratio (Fig. 10 b). Similarly to the trend observed in the domain spacing model (Fig. 9 b) but more pronounced, low solvent ratios, reflecting a mixture of toluene and THF, are associated with lower FWHM predictions at low additive ratios, indicating improved ordering from the co-solvents improving the mobility of both blocks in the BCP solution. However, at higher additive ratios, i.e., greater than 2, the relationship reverses and lower solvent ratios contribute to higher FWHM predictions. This suggests a complex relationship between the solvent processing history, the additive ratio, and the FWHM of the BCP films. Additionally, when plotting the SHAP values of additive type against the additive type itself, CN (type 0) shows tightly clustered contributions to FWHM, suggesting consistent and controlled effects on BCP film ordering (Fig. 10 c). In contrast, MN (type 1) is associated with broader and higher FWHM predictions, indicating more variability and poorer control. These results suggest that small changes in volatility of the additive could lead to measurable changes in the ordering of the BCP films.
@@ 4 Conclusion
In summary, a framework enabling the high-throughput analysis of BCP thin film morphology was developed, facilitating analysis of GISAXS and AFM characterization data. The process for extracting domain spacing from both GISAXS and AFM measurements is both highly accurate, time efficient, and provide higher resolution values than the raw data provides. Average fitting $R^{2}$ values of 0.99 and 0.96 were achieved for the GISAXS and AFM data, respectively, taking less than one second per sample. Furthermore, two-dimensional values of grain size were obtained by utilizing the AFM image analysis approach of classification, segmentation, and color-wheel or Voronoi analysis depending on the morphology. The CNN classifier proved to be highly accurate and computational inexpensive, while providing physically accurate predictions reflected by the SHAP analysis. Overall, this provides a platform for the rapid exploration of BCP morphological space by accelerating the rate of morphological analysis for the BCP films.
@@
Additionally, the rapid morphological analysis of the BCP films enabled the creation of a dataset for ML model training. RF models were highly capable of accurately classifying samples based on processing conditions, while a CNN achieved similar accuracy using AFM images directly as inputs. The RF models were also extremely accurate when predicting morphological properties measured from GISAXS, however, precision was limited when predicting outside the well-ordered regime due to an imbalance of data. Meanwhile, XGB and SVM models trained on AFM data were generally accurate but their predictive capabilities suffered from variability within the AFM data. SHAP analysis was employed to interpret the processing–structure relationships learned by the RF models from the GISAXS data, confirming that the model’s predictions were physically meaningful. Beyond validating the strong influence of additive ratio, SHAP also revealed interactions between processing history, specifically the solvent ratio and additive ratio, and the GISAXS-measured domain spacing and FWHM. Despite the similar boiling points of the two additives, the slightly lower boiling point of MN resulted in decreased domain spacing, increased FWHM, and more variability among samples relative to those with CN, indicating slight changes in the volatility of the additive have pronounced effects on BCP thin film morphology. The discussed framework combining high-throughput morphological analysis with interpretable ML has demonstrated highly accurate evaluation of BCP thin film morphology. Furthermore, the results not only provided a predictive tool for BCP thin film morphology based on processing conditions, but also a quantitative understanding of the processing–structure relationships.
@@
Future studies could benefit from this approach and focus on expanding the design space to include new polymer chemistries, molecular weights, and different solvent types. Additionally, the variation in AFM data could be greatly improved by collecting multiple replicates across each sample, as well as increasing the field of view of the images. This would not only decrease the variability of the data, but also increase the size of grains able to be measured, allowing for investigation of much more well-ordered samples. Moving forward, these changes should result in significant improvements in the performance of ML models trained on AFM data. However, collection of AFM images is currently one of the largest bottlenecks in this process, requiring significant experimentation time. Increasing both the number of AFM images and the size of the images would further exasperate this bottleneck. Therefore, utilizing a super-resolution approach could allow for the rapid collection of low-resolution AFM images which could be transformed into high resolution space.71 With the above-mentioned recommendations in mind, the exploration of BCP thin film morphologies could be greatly accelerated with the use of such an ML-enabled, integrated high-throughput data analysis approach.

### v2

@@ 3.3 Processing-structure relationship prediction
While the SHAP evaluations for the GISAXS-measured FWHM and domain spacing models share some similarities, key differences also emerge. Interestingly, the additive type has a much more significant mean SHAP value for the FWHM predictions, indicating the additive type has a more profound impact on BCP FWHM than domain spacing (Fig. 9 a left). However, additive ratio still maintains the largest mean SHAP value, 0.00029, followed by additive type, 0.00018, and solvent ratio, 0.000054. Similar to the local SHAP analysis for domain spacing, high additive ratios cluster at lower SHAP values, reflecting rapid ordering of the BCP, whereas lower additive ratios exhibit larger and more scattered SHAP values, indicative of weakly-ordered samples with broad FWHM. (Fig. 9 a right). Although the SHAP values for additive type show a similar trend for those related to the domain spacing model, they are far more distinctly separated in the FWHM model, indicating substantial variation in FWHM with changes of additive type. Diving deeper into the SHAP analysis for the FWHM prediction, an interesting relationship can be observed between solvent ratio and additive ratio (Fig. 9 b and Fig. S12). Similarly to the trend observed in the domain spacing model (Fig. 8 b) but more pronounced, low solvent ratios, reflecting a mixture of toluene and THF, are associated with lower FWHM predictions at low additive ratios, indicating improved ordering from the co-solvents improving the mobility of both blocks in the BCP solution. However, at higher additive ratios, i.e., greater than 2, the relationship reverses and lower solvent ratios contribute to higher FWHM predictions. This suggests a complex relationship between the solvent processing history, the additive ratio, and the FWHM of the BCP films. Additionally, when plotting the SHAP values of additive type against the additive type itself, CN (type 0) shows tightly clustered contributions to FWHM, suggesting consistent and controlled effects on BCP film ordering (Fig. 9 c). In contrast, MN (type 1) is associated with broader and higher FWHM predictions, indicating more variability and poorer control. These results suggest that small changes in volatility of the additive could lead to measurable changes in the ordering of the BCP films.
@@ 4 Conclusion
In summary, a framework enabling the high-throughput analysis of BCP thin film morphology was developed, facilitating analysis of GISAXS and AFM characterization data. The process for extracting domain spacing from both GISAXS and AFM measurements is highly accurate, time efficient, and provides higher resolution values than the raw data provides. Average fitting $R^{2}$ values of 0.99 and 0.96 were achieved for the GISAXS and AFM data, respectively, taking less than one second per sample. Additionally, higher-order peaks were observed in the GISAXS data, and the determination of the spacing ratio between the peaks was automated with an accuracy of 92%. Furthermore, two-dimensional values of grain size were obtained by utilizing the AFM image analysis approach of classification, segmentation, and color-wheel or Voronoi analysis depending on the surface feature type. The CNN classifier proved to be highly accurate and computationally inexpensive, while providing physically accurate predictions reflected by the SHAP analysis. Overall, this provides a platform for the rapid exploration of BCP morphological space by accelerating the rate of morphological analysis for the BCP films.
@@
Additionally, the rapid morphological analysis of the BCP films enabled the creation of a dataset for ML model training. SVM models were highly capable of accurately classifying sample domain orientation when given processing conditions, while a CNN achieved similar accuracy using AFM images directly as inputs. RF models were extremely accurate when predicting morphological properties measured from GISAXS, however, precision was limited when predicting below 34 nm due to an imbalance of data. Meanwhile, XGB and SVM models trained on AFM data were generally accurate but their predictive capabilities suffered from variability within the AFM data. SHAP analysis was employed to interpret the processing–-structure relationships learned by the RF models from the GISAXS data, confirming that the model’s predictions were physically meaningful. Beyond validating the strong influence of additive ratio, SHAP also revealed interactions between processing history, specifically the solvent ratio and additive ratio, and the GISAXS–measured domain spacing and FWHM. Despite the similar boiling points of the two additives, the slightly lower boiling point of MN resulted in decreased domain spacing, increased FWHM, and more variability among samples relative to those with CN, indicating slight changes in the volatility of the additive have pronounced effects on BCP thin film morphology. The discussed framework combining high-throughput morphological analysis with interpretable ML has demonstrated highly accurate evaluation of BCP thin film morphology. Furthermore, the results not only provided a predictive tool for BCP thin film morphology based on processing conditions, but also a quantitative understanding of the processing–structure relationships.
@@
Future studies could benefit from this approach and focus on expanding the design space to include new polymer chemistries, molecular weights, and different solvent types. This would enable analysis of a broader range of block copolymer morphologies, including gyroid and other complex structures, beyond the hexagonal cylinder morphology studied here. Such an expansion would strengthen the framework allowing for analysis of all BCP morphologies. Additionally, the variation in AFM data could be greatly improved by collecting multiple replicates across each sample, as well as increasing the field of view of the images. This would not only decrease the variability of the data, but also increase the size of grains able to be measured, allowing for investigation of much more well-ordered samples. Moving forward, these changes should result in significant improvements in the performance of ML models trained on AFM data. However, collection of AFM images is currently one of the largest bottlenecks in this process, requiring significant experimentation time. Increasing both the number of AFM images and the size of the images would further exasperate this bottleneck. Therefore, utilizing a super-resolution approach could allow for the rapid collection of low-resolution AFM images which could be transformed into high resolution space.chang_accelerating_2019 With the above-mentioned recommendations in mind, the exploration of BCP thin film morphologies could be greatly accelerated with the use of such an ML-enabled, integrated high-throughput data analysis approach.
@@ 5 Data Availability
6 Supporting Information
The Supporting Information is available free of charge at Our Paper Link. This file includes details of GISAXS bulk morphology analysis; validation and verification of AFM analysis methods; CNN architecture; morphological datasets derived from GISAXS and AFM; machine learning model comparisons and grid-search parameter space; the effect of training-set size on model performance; morphology predictions from all trained models; and SHAP interaction plots.